OpenStudy (anonymous):

am i right my answer is d -20 Evaluate. 1 – 3 • 7 + 4 • 2 A. –32 B. –28 C. –12 D. –20

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do the two multiplications first

OpenStudy (anonymous):

3*7=21 4*2=8

OpenStudy (anonymous):

giving you \[1-21+8\]

OpenStudy (anonymous):

28?

OpenStudy (anonymous):

or -28

OpenStudy (anonymous):

slow what is \(1-21\) ?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

yes then \(-20+8\) ?

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

got it as always, the answer is C

OpenStudy (anonymous):

thx @satellite73

OpenStudy (anonymous):

yw don't rush through these guessing go slow, step by step write it all down \[1-3\times 7+4\times 2\\ =1-21+8\\ =-20+8\\ =-12\]